# How Linguistics Learned to Stop Worrying and Love the Language Models

### v1

@@ 4.3 LMs are a proof of concept for gradient, usage-based theories of language
Under the real patterns account, we are open to the possibility that a wide variety of linguistic theories can be fruitfully combined with LMs. But the real patterns story is agnostic as to what theoretical framework will in fact be most explanatory and predictive. There is a separate question raised here, which is whether the LM era should in fact push us towards any particular theory for explaining language. We think that the success of LMs strengthens the case for existing usage-based language theories (closely related to the statistical tradition in 3.3.2) by serving as a model system for how these positions could work, and as proof of concept that they can work. By their very nature, graded usage-based theories have been difficult to formalize to the extent that one could, for example, generate or process sentences with them; language models are proofs of concept showing that the fuzzy, graded representations used in these theories can deliver a model of language.
@@
Traditional views in semantics often posit that words have discrete meanings and can be combined using systematic rules (Heim and Kratzer,, 1998). In contrast, a distributional view of meaning assumes meanings are best captured by usage patterns (Potts,, 2019; Goldberg,, 2019; Erk,, 2012; Baroni and Lenci,, 2010; Baroni et al.,, 2014; Elman, 1990a, ), an intuition now operationalized inside language models as continuous high-dimensional vectors that encode the contexts in which a word appears, resulting in a rich web of lexical relationships represented within the geometry of the vector space (Mikolov et al.,, 2013). If we think of the meaning of the word fire as being represented by its embedding in context, we find that the meaning of the word is slightly different in every context in which it occurs, but clustering together in space (campfires in a distinct cluster from a building fire even though these meanings are in some theories often lumped in as the same “meaning”; Chronis and Erk,, 2020). As such, LMs show that we might not need a strict separation between core meaning and a richly context-dependent and pragmatics-laden meaning in use. LMs give us a way of formalizing the intuition that meaning might be is gradient, fuzzy, and usage-based in a way that works in practice, not just in theory.
@@
LMs seem to be able to perform tasks that require knowledge of syntactic category and grammatical role, such as generalizing part of speech when encountering a novel token (Kim and Smolensky,, 2021). One might be then tempted to ask: what are the syntactic categories an LM uses? And we can, for instance, probe for grammatical category in models and see whether the categories that emerge correspond to ones posited by linguists (Chi et al.,, 2020; Papadimitriou et al.,, 2021)?
A key finding of this work is that, while there are natural clusters in part of speech and grammatical role, such clusters are necessarily fuzzy and gradient. For instance, Papadimitriou et al., (2021) found robust representations of subjecthood but they were graded with intransitive subjects and passive subjects less “subject-y” than the subjects of transitive active verbs. These representations could be discretized to make judgments like “BERT represents this word as a subject but not this one”, but this discretization would be lossy. This graded notion of syntactic category and grammatical role has a long history in functional linguistics (Ross,, 1972) and, as such, we think can be useful in giving us a new way to think about claims that word classes are not stable across languages (Haspelmath,, 2012) and that grammatical categories exist in graded hierarchies (Comrie,, 1989; Keenan and Comrie,, 1977).
@@
While traditional theories have emphasized the compositional nature of language, the construction-based approach has emphasized that not all (or even most) language can be explained strictly compositionally (Goldberg,, 2009). This position argues for a graded notion of compositionality: knowing what “green tea” is requires more than just knowing the meaning of “green” and the meaning of “tea”—even though this isn’t strictly an idiom in that the meanings of “green” and “tea” are still relevant. There has been a lot of work on whether LMs can handle strict compositionality (Lake and Baroni,, 2018; Kim and Linzen,, 2020; Russin et al.,, 2024), as well as work on how LMs handle more idiosyncratic constructions (Mahowald,, 2023; Misra and Mahowald,, 2024; Weissweiler et al., 2023a, ; Tseng et al.,, 2022; Devlin et al.,, 2019). Neural models can combine the best of both worlds in that they can maintain multiple levels of representation at once (Baroni et al.,, 2014). They don’t have to decide whether “green tea” is fully composed or fully stored: it can be a little of both. They can even be used to measure the degree of compositionality quantitatively (Rathi et al.,, 2021; Socolof et al.,, 2022).
@@
Traditional linguistic theory has often posited a cognitive separation in processing between linguistic layers (phonology, syntax, semantics, pragmatics). But evidence from psycholinguistics (e.g., Tanenhaus et al.,, 1995; Shain et al.,, 2024; Fedorenko et al.,, 2020) suggests these layers are not processed in neat discrete stages but are richly integrated. LMs do seem to have some separation such that lower layers of the model handle morphology with higher layers handling syntax and semantics (Tenney et al.,, 2019). But the nature of their architecture is to share information in hard-to-untangle nonlinear ways. So their functional separation is not strict but fuzzy and gradient—similar to psycholinguistic evidence for humans.
@@ 4.4 What the inductive biases of LMs can tell us about language
Since neural language models do have inductive biases, perhaps one reason for their success in learning language is because those inductive biases line up with those of humans. Indeed, a number of recent studies have demonstrated an apparent alignment of LMs’ inductive biases and the structure of human language.
In one such study, Kallini et al., (2024) refute the claim that neural LMs can learn any language, including unnatural ones, equally well (Bolhuis et al.,, 2024). They compare the learning curves for the GPT-2 architecture trained on language modeling on English text against models trained on various transformations of the English text, designed to create languages that are intuitively ‘impossible’. For example, one transformation applies a deterministic shuffling function to the tokens of English text, creating extraordinarily complex but deterministic word order rules that violate all known formal characterizations of syntax, and another transformation creates a new agreement marker that must appear 4 tokens away from a verb, also an unnatural pattern. Kallini et al., (2024) find that the model learns from real English text consistently faster than the baselines.
These results and others show that Transformers and related models do have inductive biases that align with human language in important ways. However, LM inductive biases don’t arise from restrictedness to a particular formal language class, as might be expected from the generative paradigm in linguistics. In fact, in terms of formal expressivity, it seems that Transformers are mismatched with the usual formal language classes used to characterize language. Whereas human language is sometimes characterized using (extensions of) the Chomsky–Schützenberger hierarchy (Chomsky and Schützenberger,, 1963; Weir,, 1988), which encompasses well-known classes such as regular and context-free languages, modern Transformers as they are currently applied (and other recently successful language model architectures such as State Space Models: Gu and Dao,, 2023) seem to inhabit formal language classes defined by circuit complexity (Merrill et al., 2022a, ; Strobl et al.,, 2024; Merrill et al., 2024a, ), a formal language hierarchy which is orthogonal to the Chomsky–Schützenberger hierarchy. [^11] Interestingly, circuit complexity has also been used to characterize the computational capacity of biologically realistic populations of neurons (Maass,, 1997; Maass and Markram,, 2004). So human performance may also be ultimately limited in this way. To the extent that Transformers have inductive biases that help (or hinder) their language learning, they likely arise from something other than the expressive limits of their architecture.
Below we consider some apparent learning biases of modern LMs which may be aligned with the structure of human language.
@@
One apparent inductive bias of autoregressive language models that seems to support language learning is information locality: a bias towards sequences where elements that statistically predict each other are close to each other. For example, in a phrase such as big brown box, the noun box and the adjective brown are highly predictive of each other—boxes, especially cardboard ones, are often brown, for many reasons—and so these words are likely to be close to each other: the alternate order brown big box sounds odd or like it is conveying some other special meaning (Futrell,, 2019; Culbertson et al.,, 2020; Scontras,, 2023; Dyer et al.,, 2023). Zooming out, paragraphs of text typically contain information about a certain topic locally, so words within a paragraph are likely to be more correlated than words spanning paragraphs. Indeed, large-scale studies of text have shown that statistical correlations among words fall off with distance (Li,, 1991; Lin and Tegmark,, 2017; Tanaka-Ishii,, 2021). Zooming in, most of the phonotactic restrictions on sound co-occurrences within words are local in nature (Heinz,, 2018).
@@
Locality is a general property of human language (Behaghel,, 1930; Givón,, 1991): words are usually contiguous units, usually modified by prefixes and suffixes (directly adjacent to them, ordered by ‘relevance’ to the root: Bybee,, 1985; Saldana et al.,, 2024), and words linked by syntactic dependencies tend to be close to each other (Gibson,, 1991, 1998; Liu,, 2008; Liu et al.,, 2017), more than would expected under random grammars within a linguistically realistic formalism (Gildea and Temperley,, 2007; Park and Levy,, 2009; Gildea and Temperley,, 2010; Futrell et al., 2015a, ; Futrell et al., 2020b, ). These locality properties may be special cases of information locality (Futrell,, 2019; Mansfield,, 2021; Hahn et al., 2021a, ; Mansfield and Kemp,, 2023). Several of the experiments of Kallini et al., (2024) (discussed above) can be interpreted as demonstrating the information locality bias in GPT-2, because they involve baseline languages which disrupt information locality.
@@
The bias towards locality seems to come from next-token prediction task performed by LMs. Thus, it is likely that any model with a similar task and loss function would have a similar bias. We believe that the bias toward information locality is an ‘ember of autoregression’ in the terminology of McCoy et al., (2023), one which may be shared with humans.
Relative insensitivity to word order
Transformers have a bias toward learning permutation-symmetric functions—that is, they struggle with learning functions that depend closely on the exact order of tokens within a sequence (Lavie et al.,, 2024). The effect of this bias has been observed empirically in Transformer-based language models, which early work found to be in many cases insensitive to word order (Sinha et al.,, 2021; Hessel and Schofield,, 2021; Pham et al.,, 2021; Gupta et al.,, 2021; O’Connor and Andreas,, 2021). Interestingly, humans also have difficulty remembering the exact ordering of long sequences of symbols, especially in the middle of a sequence (Henson,, 1998).
[…]
Learnability beyond inductive bias
The success of large language models invites us to go beyond thinking about learnability purely in terms of inductive bias, and to think about the limits of what any model or organism could learn, and what properties language should have in order to facilitate that learning. Addressing these issues has spurred serious new inquiry into fundamental questions about the relationship between statistical and linguistic structure, and what and how much it is possible to learn from a (large) corpus of language.
For example: Merrill et al., 2021a , drawing on Harris, (1954), state formal conditions under which (an emulation of) semantics may be acquired on the basis of distributional structure by any model, the key factor being to what extent semantics are expressed in ways that are not context-dependent—a factor deeply related to the systematic, compositional structure of human language. Merrill et al., 2022b furthermore show that entailment relations—logical relationships between propositions—can be extracted in principle from distributional statistics of text under certain assumptions about humans’ pragmatic behavior during language production, to wit that humans tend not to state the same fact more than once. The argument is formalized using the Rational Speech Acts model of pragmatics, a probabilistic game-theoretic model of language production and comprehension (Goodman and Frank,, 2016; Scontras et al.,, 2017; Degen,, 2023). Subsequently, in empirical work inspired by their theory, Merrill et al., 2024b find that entailment relations can be extracted from distributional statistics, but only under an inversion of their earlier theory—it was in fact necessary to assume a high level of redundancy in text, revising the earlier assumption.
This is just one example of a sequence of scientific models and studies, inspired by and focusing on language models, which have brought real new insight about the relationship between the structure of language (in this case, semantics), the logic of language use in context (pragmatics), and the information-theoretic structure of large bodies of language as produced by humans (distributional structure). The back-and-forth of theory development, testing against data, and theory revision in response to the data is the hallmark of healthy science.
@@ 4.5 Functional explanations for human language
Another route to explanatory adequacy avoids appeals to innateness and language-specific learning biases. The reason that ‘unnatural’ languages don’t occur may be something other than their putative innate non-learnability. Unnatural languages may not exist because they would be a poor fit to the function of language: communication and social coordination under general cognitive constraints on how language may be produced and comprehended (Chomsky,, 2005; Gibson et al.,, 2019; Levshina,, 2022; Bickel et al.,, 2024). Unsurprisingly, this function-based approach to explaining the structure of human language is associated with the functionalist school of linguistics, which is often contrasted to the generative or formalist approach (Newmeyer,, 1998), and which often overlaps with the statistical tradition discussed in Section 3.3.2. In terms of the competence–performance distinction, functionalist approaches hold that the structure of language ultimately arises from constraints and pressures on performance, which over time shape the grammars of languages that we see (Hawkins,, 1994, 2004, 2014; Haspelmath,, 2008; Comrie,, 1989).
On some level, functional explanations of human language are more basic than explanations based on learning biases. Even if we have an innate inductive bias toward human languages, this does not answer the ultimate question of why language is the way it is, it just moves the question back to why humans have one inductive bias rather than another. [^12] Indeed, a number of human learning biases tested in the artificial language learning experiments are those that also increase the communicative efficiency of language (for example, Fedzechkina et al.,, 2012; Culbertson and Adger,, 2014; Fedzechkina et al.,, 2017). From the functionalist perspective, it is no surprise that the inductive biases observed in Transformers above seem to be aligned with properties of language in key ways. They both may have their basis in deeper information-processing constraints.
One of the major findings in language comprehension has been that, like much of human cognition (Bar,, 2009; Friston and Kiebel,, 2009), it is largely predictive in nature (Ryskin and Nieuwland,, 2023): the way that processing works is influenced by the predictability of each word (or sub-word unit) conditional on previous words (Hale,, 2001; Levy,, 2008; Wilcox et al.,, 2020; Wilcox et al., 2023b, ; Xu et al.,, 2023). Furthermore, language comprehension is highly incremental (Tanenhaus et al.,, 1995; Smith and Levy,, 2013), meaning that information is processed and integrated as quickly as possible as it comes in, largely without buffering or apparent stages. There is also a high degree of incrementality in language production (Ferreira and Swets,, 2002), and incremental predictability plays a key role there too: for example, disfluencies and pauses in speech happen before unpredictable words (Goldman-Eisler,, 1957; Beattie and Butterworth,, 1979; Dammalapati et al.,, 2019), and unpredictable words are enunciated more slowly (Bell et al.,, 2009). Experimental psycholinguistics has delivered a picture of human language processing that is deeply entwined with the task of autoregressive prediction, the same task that machine learning researchers have converged on for LM training.
@@
In general, we believe that the functional level of explanation is where linguistics and language modeling make the most fruitful contact. While bird wings and airplane wings are very different things, they share the function of flying in the Earth’s atmosphere, and so the engineering principles of aerodynamics are relevant to both airplane engineers and ornithologists (Marr,, 1982). Similarly, neural language models and human language are different things, but they share relevant principles at the level of function: they both encode information and decode information incrementally and predictively. To the extent that human language is structured to support this kind of incremental prediction (for example, through information locality), it is also aligned with the strengths and inductive biases of language models.
In fact, the explosive growth of language models and neural architectures for language provides a unique and exciting opportunity to develop the functionalist approach to language, in which models of language processing and models of language structure are like two sides of a coin (Jaeger and Tily,, 2011; Gibson et al.,, 2019; Futrell and Hahn,, 2022).
On the side of developing models of human language processing, neural language models have become an indispensable tool. Even after converging on the idea that predictive models were important for language processing, psycholinguistics as a field was hampered by the fact that strong probabilistic models for language were not available, resulting in a number of conclusions drawn on the basis of what now seem to be fairly weak models (Frank and Bod,, 2011; Fossum and Levy,, 2012; Smith and Levy,, 2013). Now, neural models make it possible to probe the limits of predictive models of language processing (Wilcox et al.,, 2020; Wilcox et al., 2023c, ; Xu et al.,, 2023), with the discovery that certain phenomena seem to go beyond the simplest predictive models (van Schijndel and Linzen,, 2018, 2021; Huang et al.,, 2024; Staub,, 2024), and that human processing is better characterized by prediction under resource constraints similar to but more severe than those present in neural models (Futrell et al., 2020a, ; Hahn et al.,, 2022; Kuribayashi et al.,, 2022; Timkey and Linzen,, 2023; Oh and Schuler,, 2023; De Varda and Marelli,, 2024).
@@
On the side of explaining language structure, language models have again proved a key tool and source of inspiration. A sizable literature exists about what kinds of languages emerge in simulated populations of agents who communicate by encoding and decoding meanings into strings using various neural architectures, and what constraints (on the environment or the agents) are necessary for the emergent languages to resemble natural language (Lazaridou et al.,, 2017; Mordatch and Abbeel,, 2018; Steinert-Threlkeld,, 2020; Kuciński et al.,, 2021). Hahn et al., (2020) show that certain universal properties of word order, identified by Greenberg, 1963b (see also Dryer,, 1992), can be derived by finding grammars that optimize an objective function defined by (1) how easily a parse tree can be recovered from a string, and (2) how much surprisal is incurred by each word in context, with both terms operationalized using neural-network models. Because human language processing appears to be probabilistic and predictive, any theory of linguistic structure based on processing efficiency must be evaluated using a strong probabilistic predictive model, and language models provide exactly this tool. Overall, we think it is not a coincidence that the function of language models coincides so closely with the proposed function of natural language. Much of the structure of language seems to support a kind of local incremental predictability that dovetails with the function of language models (Futrell and Hahn,, 2022, 2024), and we believe that this property of language drives much of the success of autoregressive LMs.

### v2

@@ 4.3 LMs are a proof of concept for gradient, usage-based theories of language
The real patterns account leaves open the possibility that a wide variety of linguistic theories can be fruitfully combined with LMs. But there is a separate question raised here, which is whether the LM era should push us towards any particular theory for explaining language. We believe that the success of LMs strengthens the case for existing usage-based language theories based on gradient representations. By their very nature, such theories have been difficult to formalize to the extent that one could, for example, generate or process sentences with them. LMs serve as model systems for how these ideas can be implemented, and as proof of concept that they can work.
@@
Traditional views in semantics often posit that words have discrete meanings and can be combined using systematic rules (Heim and Kratzer,, 1998). In contrast, a distributional view of meaning assumes meanings are best captured by usage patterns (Potts,, 2019; Goldberg,, 2019; Erk,, 2012; Baroni and Lenci,, 2010; Baroni et al.,, 2014; Elman, 1990a, ), an intuition now operationalized inside language models as continuous high-dimensional vectors that encode the contexts in which a word appears, resulting in a rich web of lexical relationships represented within the geometry of the word vector space (Mikolov et al.,, 2013). If we think of the meaning of the word fire as being represented by its embedding in context, we find that the meaning of the word is slightly different in every context in which it occurs, but clustering together in space (campfires in a distinct cluster from a building fire even though these meanings are in some theories often lumped in as the same “meaning”; Chronis and Erk,, 2020). As such, LMs show that we might not need a strict separation between core meaning and a richly context-dependent and pragmatics-laden meaning in use. LMs give us a way of formalizing the intuition that meaning might be gradient, fuzzy, and usage-based in a way that works in practice, not just in theory.
@@
LMs seem to be able to perform tasks that require knowledge of syntactic category and grammatical role, such as generalizing part of speech when encountering a novel token (Kim and Smolensky,, 2021). One might then be tempted to ask: what are the syntactic categories an LM uses? And we can, for instance, probe for grammatical category in models and see whether the categories that emerge correspond to ones posited by linguists (Chi et al.,, 2020; Papadimitriou et al.,, 2021)? A key finding of this work is that, while there are natural clusters in part of speech and grammatical role, such clusters are necessarily fuzzy and gradient. For instance, Papadimitriou et al., (2021) found robust representations of subjecthood. But they were graded, with intransitive subjects and passive subjects less “subject-y” than the subjects of transitive active verbs. These representations could be discretized to make judgments like “BERT represents this word as a subject but not this one”, but this discretization would be lossy. This graded notion of syntactic category and grammatical role has a long history in functional linguistics (Ross,, 1972) and, as such, can be useful in giving us a new way to think about claims that word classes are not stable across languages (Haspelmath,, 2012) and that grammatical categories exist in graded hierarchies (Comrie,, 1989; Keenan and Comrie,, 1977).
@@
While traditional theories have emphasized the compositional nature of language, the construction-based approach has emphasized that not all (or even most) language can be explained strictly compositionally (Goldberg,, 2009). This position argues for a graded notion of compositionality: knowing what “green tea” is requires more than just knowing the meaning of “green” and the meaning of “tea”—even though this isn’t strictly an idiom in that the meanings of “green” and “tea” are still relevant. There has been a lot of work on whether LMs can handle strict compositionality (Lake and Baroni,, 2018; Kim and Linzen,, 2020; Russin et al.,, 2024), as well as work on how LMs handle more idiosyncratic constructions (Mahowald,, 2023; Misra and Mahowald,, 2024; Weissweiler et al.,, 2023; Tseng et al.,, 2022; Devlin et al.,, 2019). Neural models can combine the best of both worlds in that they can maintain multiple levels of representation at once (Baroni et al.,, 2014). They don’t have to decide whether “green tea” is fully composed or fully stored: it can be a little of both. They can even be used to measure the degree of compositionality quantitatively (Rathi et al.,, 2021; Socolof et al.,, 2022).
@@
Traditional linguistic theory has often posited a cognitive separation in processing between linguistic layers (phonology, syntax, semantics, pragmatics). But evidence from psycholinguistics (e.g., Tanenhaus et al.,, 1995; Shain et al.,, 2024; Fedorenko et al.,, 2020) suggests these layers are not processed in neat discrete stages but are richly integrated. LMs do seem to have some separation such that, for example, lower layers of the model handle morphology with higher layers handling syntax and semantics (Tenney et al.,, 2019). However, the nature of their architecture is to share information in hard-to-untangle nonlinear ways. So their functional separation is not strict but fuzzy and gradient—similar to psycholinguistic evidence for humans.
@@ 4.4 What the inductive biases of LMs can tell us about language
Since neural language models do have inductive biases, perhaps one reason for their success in learning language is that those inductive biases meaningfully align with linguistic structure. Indeed, a number of recent studies have investigated this possibility.
In one such study, Kallini et al., (2024) refute the claim that neural LMs can learn any language, including unnatural ones, equally well (Bolhuis et al.,, 2024). They compare the learning curves for the GPT-2 architecture trained on language modeling on English text against models trained on various transformations of the English text, designed to create languages that are intuitively ‘impossible’, but which still have the same level of overall predictability as the original English text. For example, one transformation applies a deterministic shuffling function to the tokens of English text, creating extraordinarily complex but deterministic word order rules that violate all known formal characterizations of syntax, and another transformation creates a new agreement marker that must appear exactly 4 tokens away from a verb, also an unnatural pattern. Kallini et al., (2024) find that the model learns from real English text consistently faster than these baselines (see also Mitchell and Bowers,, 2020; Yang et al.,, 2025; Xu et al.,, 2025; Ziv et al.,, 2025, among others).
These results and others show that Transformers and related models have inductive biases that align with human language. However, the major determinant of inductive biases in LM is not that they are restricted to a particular formal language class, as might be expected from the generative linguistics paradigm. In fact, in terms of formal expressivity, it seems that Transformers are mismatched with the usual formal language classes used to characterize language. Whereas human language is sometimes characterized using (extensions of) the Chomsky–Schützenberger hierarchy (Chomsky and Schützenberger,, 1963; Vijay-Shanker et al.,, 1987; Weir,, 1988), which encompasses well-known classes such as regular and context-free languages, modern Transformers as they are currently applied (and other recently successful language model architectures such as State Space Models: Gu and Dao,, 2024) seem to inhabit formal language classes defined by circuit complexity (Merrill et al.,, 2022; Strobl et al.,, 2024; Merrill et al.,, 2024), a formal language hierarchy which is orthogonal to the Chomsky–Schützenberger hierarchy. [^14] Interestingly, circuit complexity has also been used to characterize the computational capacity of biologically realistic populations of neurons (Maass,, 1997; Maass and Markram,, 2004). So human performance may also be ultimately limited in this way. To the extent that Transformers have inductive biases that help (or hinder) their language learning, they likely arise from something other than the expressive limits of their architecture.
Below we consider two apparent learning biases of modern LMs which may be aligned with the structure of human language: information locality and low sensitivity.
@@
Human languages are structured in a way such that elements that statistically predict each other are usually close to each other. For example, in phrase such as big brown box, the noun box and the adjective brown are highly predictive of each other—boxes, especially cardboard ones, are often brown, for many reasons—and so these words are likely to be close to each other: the alternate order brown big box sounds odd or like it is conveying some other special meaning (Futrell,, 2019; Culbertson et al.,, 2020; Scontras,, 2023; Dyer et al.,, 2023). Locality ideas of this kind pervade human language (Behaghel,, 1930; Givón,, 1991; Futrell,, 2019; Mansfield,, 2021; Hahn et al., 2021a, ; Mansfield and Kemp,, 2023): words are usually contiguous units, usually modified by prefixes and suffixes (directly adjacent to them, ordered by ‘relevance’ to the root: Bybee,, 1985; Saldana et al.,, 2024), and words linked by syntactic dependencies tend to be close to each other (Gibson,, 1991, 1998; Liu,, 2008; Liu et al.,, 2017), more than would expected under random grammars within a linguistically realistic formalism (Gildea and Temperley,, 2007; Park and Levy,, 2009; Gildea and Temperley,, 2010; Futrell et al.,, 2015; Futrell et al., 2020b, ).
@@
Autoregressive LMs such as GPT-2 also show a bias towards information locality, as demonstrated by several of the experiments of Kallini et al., (2024) (discussed above): many of the counterfactual languages which are harder to learn in those experiments are also those that disrupt information locality. The bias towards locality seems to come from the next-token prediction task performed by LMs. The bias toward information locality is an ‘ember of autoregression’ in the terminology of McCoy et al., (2023), one which helps language learning and is likely shared with humans.
Relatively low sensitivity
Another related inductive bias in Transformers is the bias toward learning functions with low sensitivity or low polynomial degree (Hahn et al., 2021b, ; Abbe et al.,, 2023; Bhattamishra et al.,, 2023). Sensitive functions are functions on input strings whose outputs change drastically based on small changes to the input. For example, a function on input bitstrings that counts the parity of the input—returning odd or even depending on the number of 1’s in the input—is maximally sensitive and high-degree, because any single change to the input bits will flip the output of the function (O’Donnell,, 2014). Although the Transformer architecture has the ability to represent highly sensitive functions in terms of its representational capacity, this turns out not to be relevant to its inductive bias in learning. Rather, the bias toward low-sensitivity functions comes from the shape of the loss landscape induced by the model. In particular, any parameter setting representing a highly sensitive function in the Transformer architecture must be brittle, meaning that a small change to the parameters would make the Transformer produce some different, lower-sensitivity function (Hahn and Rofin,, 2024). Thus high-sensitivity functions are unlikely to be reached through a gradient-descent-based learning process.
Human languages, viewed (for example) as functions from strings to meanings or to grammaticality judgments, also seem to be relatively low-sensitivity (Hahn et al., 2021b, ). We do not find human languages where, for example, a word is well-formed if and only if it contains an odd number of some certain segment, even though it would be easy to express this language using a regular grammar.
However, like information locality, low sensitivity is not an absolute formal restriction on languages. For example, calculating the meaning of iterated negation is like a parity function: What does it mean when someone says they are not not not not not not wearing a tie? These high-sensitivity parts of language are nonetheless rare in usage, and difficult to understand in practice. Relative low sensitivity is an important statistical property of human language, perhaps reflecting a general cognitive constraint for humans, which has come to light as the result of taking the inductive biases of LM architectures seriously.
4.5 Convergence of LMs and psycholinguistics on predictive processing
It is not a coincidence that the basic paradigm for training language models—incremental prediction of upcoming input—mirrors ideas about how human language processing works. In early neural network work, Elman, 1990b cites two reasons for choosing to use a prediction task in training his models: “minimal assumptions about special knowledge required for training” and the fact that “much of what listeners do involves anticipation of future input”. Indeed, there is a deep functional convergence between the predictive approach to language modeling and the ways that human language is processed and structured.
Elman, 1990a’s statement about what listeners do is not speculative. Experimental psycholinguistics has delivered a picture of human language processing that is deeply entwined with the task of autoregressive prediction, mirroring other areas of cognition (Rao and Ballard,, 1999; Bar,, 2009; Friston and Kiebel,, 2009; Ryskin and Nieuwland,, 2023). In comprehension (the process of decoding meaning from linguistic form in real time), the way that human processing works is influenced by the predictability of each word (or sub-word unit) conditional on previous words (Hale,, 2001; Levy,, 2008; Wilcox et al.,, 2020; Wilcox et al., 2023b, ; Xu et al.,, 2023). Furthermore, language comprehension is highly incremental (Tanenhaus et al.,, 1995; Smith and Levy,, 2013), meaning that information of different kinds is processed and integrated as quickly as possible as it comes in. There is also a high degree of incrementality in language production (Ferreira and Swets,, 2002), and incremental predictability plays a key role there too: for example, disfluencies and pauses in speech happen before unpredictable words (Goldman-Eisler,, 1957; Beattie and Butterworth,, 1979; Dammalapati et al.,, 2019; Futrell,, 2023), and unpredictable words are enunciated more slowly (Bell et al.,, 2009; Upadhye and Futrell,, 2025).
Many of these studies about the effects of predictability in language processing would have been impossible without large language models. Even after converging on the idea that prediction is an important part of language processing, psycholinguistics as a field was hampered by the fact that strong probabilistic models for language were not available, resulting in conclusions based on what now seem to be fairly weak models and methods of estimating incremental probability (Frank and Bod,, 2011; Fossum and Levy,, 2012; Smith and Levy,, 2013). Now, neural models make it possible to probe the limits of predictive models of language processing (Wilcox et al.,, 2020; Wilcox et al., 2023b, ; Xu et al.,, 2023), including the discovery that certain phenomena seem to go beyond the simplest predictive models (van Schijndel and Linzen,, 2018, 2021; Huang et al.,, 2024; Staub,, 2024), and that human processing is better characterized by prediction under resource constraints similar to but more severe than those present in neural models (Futrell et al., 2020a, ; Hahn et al.,, 2022; Kuribayashi et al.,, 2022; Timkey and Linzen,, 2023; Oh and Schuler,, 2023; De Varda and Marelli,, 2024; Clark et al.,, 2025).
@@ 4.6 Functional explanations for human language
LMs also orient us toward another route to explanatory adequacy in linguistic theory, one which posits that the form of language is related to its function: communication of thought and social coordination under general cognitive constraints on how language is produced and comprehended (Chomsky,, 2005; Gibson et al.,, 2019; Levshina,, 2022; Bickel et al.,, 2024). The functionalist school of linguistics, which is often contrasted to the generative or formalist approach (Newmeyer,, 1998), holds that the structure of language ultimately reflects constraints and pressures arising from the way language is actually used (Hawkins,, 1994, 2004, 2014; Haspelmath,, 2008; Comrie,, 1989)—perhaps motivated by factors such as a language’s “niche” such that languages with larger communities of speakers or more second language learners might have different pressures (Lupyan and Dale,, 2010; Raviv et al.,, 2019).
@@
We believe that linguistics and language modeling can make (and in fact already have made) fruitful contact at this functional level of explanation. By analogy, bird wings and airplane wings are very different things, but they share the function of flying in the Earth’s atmosphere, and so they are both shaped by the constraints of aerodynamics (Marr,, 1982; Gill,, 1995). Similarly, neural language models and human language processing mechanisms are different things, but at the level of function, they both encode and decode information in linguistic strings incrementally and predictively. To the extent that human language is structured in a way that supports this kind of incremental prediction (for example, through information locality), it is also aligned with the strengths and inductive biases of language models.
@@
Indeed, language models have already proved a key tool in functionalist models of why languages are the way they are. A sizable literature exists about what kinds of languages emerge in simulated populations of agents who communicate by encoding and decoding meanings into strings using various neural architectures, and what constraints (on the environment, the agents, or the task) are necessary for the emergent languages to resemble natural language (Lazaridou et al.,, 2017; Mordatch and Abbeel,, 2018; Steinert-Threlkeld,, 2020; Kuciński et al.,, 2021; Chaabouni et al.,, 2021). For example, Hahn et al., (2020) show that certain universal properties of word order (Greenberg,, 1963; Dryer,, 1992) can be derived by finding grammars that optimize (1) the ease of recovering a parse tree from a string, and (2) the ease of incrementally predicting each word, with both factors operationalized using neural networks (see also Kuribayashi et al.,, 2024). Clark et al., (2023) show that natural language word order seems to be structured in a way that minimizes the variance of word-by-word surprisals, again measured using neural LMs, in keeping with the theory of Uniform Information Density (Fenk and Fenk,, 1980; Levy and Jaeger,, 2007; Jaeger,, 2010; Jaeger and Tily,, 2011). More generally, because human language processing is highly probabilistic and predictive, theories of linguistic structure based on functional constraints must be evaluated using a strong probabilistic predictive model. Language models provide exactly that.